# Homogeneous solutions to the Einstein-matter equations with a magnetic field and a conformal gauge singularity

### v1

@@ 1 Introduction
This article is an extension of [15] to other matter models. We work with the same symmetry class, along with a homogeneous magnetic field satisfying the source-free Maxwell equations, and either a radiation fluid (that is, a perfect fluid with pressure equal to one-third of density) or a kinetic theory source specified by a distribution function $f$ and either collisionless (the Vlasov case) or with collisions and subject to the Boltzmann equation. This symmetry class fits very well with a homogeneous magnetic field and this combined with other fields has been much studied [5, 10, 11, 12, 16], though not, as here, with an initial conformal gauge singularity. Our method leans heavily on the existence and uniqueness theorem for Fuchsian systems, Theorem 4 of [15], which is a strengthening of the earlier result of [19]. In the different cases, we first conformally rescale the metric and matter fields and then choose variables so as to cast the Einstein equations into Fuchsian form and, in the relevant case, to regularise the Boltzmann equation. The Fuchsian system imposes certain conditions, conveniently also called Fuchsian conditions, on the initial data so it is necessary to show that there exist initial data satisfying these conditions. Then Theorem 4 of [15] requires differentiability of the coefficients of the Fuchsian system, which must be verified, and requires conditions on the eigenvalues of a matrix defined by the Fuchsian system, one condition sufficient for the existence of a solution bounded at time zero, i.e. near the (removed) singularity, and a stronger one sufficient for existence of a solution differentiable at time zero. In the case of the radiation fluid and magnetic field, with or without cosmological constant, the Fuchsian conditions are seen to be so strong that there do not exist initial data with a nonzero magnetic field. After observing this, therefore, we drop the perfect fluid case and consider the two kinetic theory cases. Both of these have a satisfactory solution: given any distribution function compactly supported in momentum space away from the origin we obtain initial data to the rescaled Einstein-Vlasov and Einstein-Boltzmann system. The precise statement can be found in Proposition 1 and Proposition 2 respectively. Based on this we prove the existence and uniqueness of solutions to the rescaled Einstein-Vlasov and Einstein-Boltzmann system. If the magnetic field is small in a certain sense which will be specified later, we even obtain that the solutions to the Einstein equations are differentiable in the rescaled, i.e. unphysical, coordinates. The precise statements are Theorem 1 and Theorem 3 respectively. This implies in particular that in the absence of a magnetic field we obtain a stronger result in the Vlasov case that was obtained previously in [4]. The reason for this is that we apply Theorem 4 of [15] which is stronger than the theorem used in [4] which is based on [19].
@@
The results concerning the rescaled Einstein-Vlasov and rescaled Einstein-Boltzmann system are also expressed in terms of the unrescaled, physical quantities. In particular we obtain for both cases asymptotic expansions close to the initial conformal gauge singularity which can be found in Theorem 2 and Theorem 4 respectively. As a result of the differentiability of the solutions for small magnetic fields we obtain sharper asymptotics in that cases. Note that we have allowed in all cases the existence of a cosmological constant, which however does not play an important role in the arguments.
@@
The paper is organised as follows. In the next section we introduce all the relevant equations where we couple the Einstein equations with a radiation fluid, collisionless matter and with matter which interacts via the Boltzmann equation for a certain range of soft potentials. In Section 3 introducing certain time variable changes we obtain the conformally rescaled equations for all the cases coming to the conclusion that in the presence of a magnetic field there is no solution to the Einstein equations coupled to a radiation fluid and a conformal gauge singularity. In Section 4 we consider the Vlasov case and obtain the main results for that case proving that all conditions to apply Theorem 4 of [15] are satisfied. Using the results concerning the Boltzmann equation of [15] and the results developed in Section 4 we obtain the main results concerning the Boltzmann case in the last section.
@@ 2.1 The metric and curvature
$=\frac{1}{2}\frac{d\tilde{k}}{dt}+\frac{1}{4}\tilde{k}^{mn}\tilde{k}_{mn}+\frac{1}{2}\tilde{b}^{ij}\frac{d}{dt}\tilde{k}_{ij}+\frac{1}{4}\tilde{k}^{2}-\frac{1}{2}\tilde{k}_{im}\tilde{b}^{ij}\tilde{b}^{mn}\tilde{k}_{nj}.$ (10)
@@
$\frac{d}{dt}\tilde{k}=\frac{d}{dt}\left(\tilde{b}^{ij}\tilde{k}_{ij}\right)=-\tilde{k}^{ij}\tilde{k}_{ij}+\tilde{b}^{ij}\frac{d}{dt}\tilde{k}_{ij},$ (11)
@@
$\tilde{R}=\frac{d\tilde{k}}{dt}+\frac{1}{4}\tilde{k}^{2}+\frac{1}{4}\tilde{k}^{ij}\tilde{k}_{ij}.$ (12)
@@
$\tilde{G}_{00}=\tilde{R}_{00}-\frac{1}{2}\tilde{g}_{00}\tilde{R}=\frac{1}{8}\tilde{k}^{2}-\frac{1}{8}\tilde{k}^{ij}\tilde{k}_{ij}.$ (13)
@@ 2.2.1 The radiation perfect fluid
The first, which is instructive but turns out to have no interesting solutions, is the radiation perfect fluid, for which the stress tensor is
@@ 2.2.2 The source-free magnetic field
$\tilde{\eta}_{ijk}=(\tilde{a})^{-1/2}\epsilon_{ijk},$ (22)
@@ 2.2.3 The kinetic part
$\tilde{T}^{\mbox{kt}}_{00}$ $=$ $\frac{1}{\sqrt{\tilde{a}}}\int_{N_{x}}f(\tilde{b}^{mn}p_{m}p_{n})^{1/2}d^{3}p,$ (35)
$\tilde{T}^{\mbox{kt}}_{0i}$ $=$ $-\frac{1}{\sqrt{\tilde{a}}}\int_{N_{x}}fp_{i}d^{3}p,$ (36)
$\tilde{T}^{\mbox{kt}}_{ij}$ $=$ $\frac{1}{\sqrt{\tilde{a}}}\int_{N_{x}}fp_{i}p_{j}(\tilde{b}^{mn}p_{m}p_{n})^{-1/2}d^{3}p,$ (37)
@@ 3.1 Behaviour of the metric and curvature
From (7,8,9) the physical Ricci tensor can be written
@@ 3.2 Behaviour of the source-terms
$\tilde{T}^{\mbox{kt}}_{00}$ $=$ $\tau^{-4}\frac{1}{\sqrt{a}}\int_{N_{x}}f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p,$ (56)
$\tilde{T}^{\mbox{kt}}_{0i}$ $=$ $-\tau^{-3}\frac{1}{\sqrt{a}}\int_{N_{x}}fp_{i}d^{3}p,$ (57)
$\tilde{T}^{\mbox{kt}}_{ij}$ $=$ $\tau^{-2}\frac{1}{\sqrt{a}}\int_{N_{x}}fp_{i}p_{j}(b^{mn}p_{m}p_{n})^{-1/2}d^{3}p.$ (58)
@@ 3.3 The rescaled Einstein equations
$\frac{2}{\tau}Z_{ij},\quad Z_{ij}=\tau 8\pi T_{ij}-\frac{1}{\tau}a_{ij},$ (60)
@@
$Z_{ij}=\frac{1}{\tau}\left[\frac{8\pi}{\sqrt{a}}\int_{N_{x}}fp_{i}p_{j}\frac{d^{3}p}{(b^{mn}p_{m}p_{n})^{1/2}}+\frac{2}{a}(a_{ij}a_{mn}-2a_{im}a_{jn})h^{m}h^{n}-a_{ij}\right],$ (61)
@@
$\frac{d}{d\tau}a=ka,\quad\frac{d}{d\tau}b^{ij}=-k^{ij},$ (64)
@@ 3.4 The Einstein constraint equations
One of the constraint equations is
$\tilde{G}_{0i}-8\pi\tilde{T}_{0i}=0\iff\frac{1}{\sqrt{\tilde{a}}}\int_{N_{x}}fp_{i}d^{3}p=0,$ (69)
@@
$\frac{1}{\tau^{-3}\sqrt{{a}}}\int_{N_{x}}fp_{i}d^{3}p=0.$ (70)
This implies that if $\int_{N_{x}}fp_{i}d^{3}p=0$ initially it will remain zero.
On the other hand we have the Hamiltonian constraint as
@@
$\iff$ $\tilde{H}=\frac{1}{8}\tilde{k}^{2}-\frac{1}{8}\tilde{k}^{ij}\tilde{k}_{ij}-\Lambda-\frac{2}{\tilde{a}}\tilde{a}_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{\tilde{a}}}\int_{N_{x}}f(\tilde{b}^{mn}p_{m}p_{n})^{1/2}d^{3}p=0.$ (71)
@@
$\frac{d}{dt}\tilde{H}$ $=\frac{1}{4}\tilde{k}\frac{d}{dt}\tilde{k}-\frac{1}{8}\tilde{k}^{ij}\frac{d}{dt}\tilde{k}_{ij}+\frac{1}{8}\tilde{k}_{ij}\tilde{k}^{im}\tilde{b}^{jn}\tilde{k}_{mn}+\frac{1}{8}\tilde{k}_{ij}\tilde{b}^{im}\tilde{k}^{jn}\tilde{k}_{mn}-\frac{1}{8}\tilde{k}_{ij}\tilde{b}^{im}\tilde{b}^{jn}\frac{d}{dt}\tilde{k}_{mn}$
$+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{k}^{mn}\tilde{T}_{mn}^{\mbox{kt}}$
$=\frac{1}{2}\tilde{k}\left(-\frac{1}{4}\tilde{k}^{mn}\tilde{k}_{mn}-\tilde{R}_{00}\right)+\frac{1}{2}\tilde{k}^{ij}\left(\frac{1}{4}\tilde{k}\tilde{k}_{ij}-\frac{1}{2}\tilde{k}_{im}\tilde{b}^{mn}\tilde{k}_{nj}-\tilde{R}_{ij}\right)+\frac{1}{4}\tilde{k}_{ij}\tilde{b}^{im}\tilde{k}^{jn}\tilde{k}_{mn}$
$+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{k}^{mn}\tilde{T}_{mn}^{\mbox{kt}}$
$=$ $-\frac{1}{2}\tilde{k}\tilde{R}_{00}-\frac{1}{2}\tilde{k}^{ij}\tilde{R}_{ij}+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{k}^{mn}\tilde{T}_{mn}^{\mbox{kt}}=-\frac{1}{2}\tilde{k}\tilde{H},$ (72)
which implies that if the constraint equation is satisfied initially it will always be satisfied.
@@
$H=\frac{1}{8}\left(\frac{k}{\tau}+\frac{6}{\tau^{2}}\right)^{2}-\frac{1}{8\tau^{4}}(\tau k_{ij}+2a_{ij})(\tau k^{ij}+2b^{ij})-\Lambda-\frac{2}{a\tau^{4}}a_{ij}h^{i}h^{j}$
$-\frac{8\pi}{\tau^{4}\sqrt{a}}\int_{N_{x}}f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p=0,$ (73)
@@
$\tau^{4}H=-\Lambda\tau^{4}+\frac{1}{8}\tau^{2}\left(k^{2}-k_{ij}k^{ij}\right)+k\tau+3-\frac{2}{a}a_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{a}}\int_{N_{x}}f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p=0.$ (74)
@@
$0=3-\frac{2}{a}a_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{a}}\int_{N_{x}}f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p.$ (75)

### v2

@@ 1 Introduction
This article is an extension of [15] to other matter models. We work with the same symmetry class, along with a homogeneous magnetic field satisfying the source-free Maxwell equations, and either a radiation fluid (that is, a perfect fluid with pressure equal to one-third of density) or a kinetic theory source specified by a distribution function $f$ and either collisionless (the Vlasov case) or with collisions and subject to the Boltzmann equation. This symmetry class fits very well with a homogeneous magnetic field and this combined with other fields has been much studied [5, 10, 11, 12, 16], though not, as here, with an initial conformal gauge singularity. Our method leans heavily on the existence and uniqueness theorem for Fuchsian systems, Theorem 4 of [15], which is a strengthening of the earlier result of [19]. In the different cases, we first conformally rescale the metric and matter fields and then choose variables so as to cast the Einstein equations into Fuchsian form and, in the relevant case, to regularise the Boltzmann equation. The Fuchsian system imposes certain conditions, conveniently also called Fuchsian conditions, on the initial data so it is necessary to show that there exist initial data satisfying these conditions. Then Theorem 4 of [15] requires differentiability of the coefficients of the Fuchsian system, which must be verified, and requires conditions on the eigenvalues of a matrix defined by the Fuchsian system, one condition sufficient for the existence of a solution bounded at time zero, i.e. near the (removed) singularity, and a stronger one sufficient for existence of a solution differentiable at time zero. In the case of the radiation fluid and magnetic field, with or without cosmological constant, the Fuchsian conditions are seen to be so strong that there do not exist initial data with a nonzero magnetic field. After observing this, therefore, we drop the perfect fluid case and consider the two kinetic theory cases. Both of these have a satisfactory solution: given any distribution function compactly supported in momentum space away from the origin we obtain initial data to the rescaled Einstein-Vlasov and Einstein-Boltzmann system. The precise statement can be found in Proposition 1 and Proposition 2 respectively. Based on this we prove the existence and uniqueness of solutions to the rescaled Einstein-Vlasov and Einstein-Boltzmann system. If the magnetic field is small in a certain sense which will be specified later, we even obtain that the solutions to the Einstein equations are differentiable in the rescaled, i.e. unphysical, coordinates. The precise statements are Theorem 1 and Theorem 3 respectively. This implies in particular that in the absence of a magnetic field we obtain a stronger result in the Vlasov case than was obtained previously in [4]. The reason for this is that we apply Theorem 4 of [15] which is stronger than the theorem used in [4] which is based on [19].
@@
The results concerning the rescaled Einstein-Vlasov and rescaled Einstein-Boltzmann system are also expressed in terms of the unrescaled, physical quantities. In particular we obtain for both cases asymptotic expansions close to the initial conformal gauge singularity which can be found in Theorem 2 and Theorem 4 respectively. As a result of the differentiability of the solutions for small magnetic fields we obtain sharper asymptotics in that case. Note that we have allowed in all cases the existence of a cosmological constant, which however does not play an important role in the arguments.
@@
The paper is organised as follows. In the next section we introduce all the relevant equations where we couple the Einstein equations with a radiation fluid, collisionless matter and with matter which interacts via the Boltzmann equation for a certain range of soft potentials. In Section 3 introducing certain time variable changes we obtain the conformally rescaled equations for all the cases coming to the conclusion that in the presence of a magnetic field there is no solution to the Einstein equations coupled to a radiation fluid and a conformal gauge singularity. In the end of that section we will state our main theorems concerning solutions to the Einstein-Vlasov and the Einstein-Boltzmann system which will be proven in Section 4 and 5 respectively. In Section 4 we consider the Vlasov case and prove the main results for that case showing that all conditions to apply Theorem 4 of [15] are satisfied. Using the results concerning the Boltzmann equation of [15] and the results developed in Section 4 we prove the main results concerning the Boltzmann case in the last section.
@@ 2.1 The metric and curvature
$=\frac{1}{2}\frac{d\tilde{k}}{dt}-\frac{1}{4}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{ij}\tilde{k}_{mn}+\frac{1}{2}\tilde{b}^{ij}\frac{d}{dt}\tilde{k}_{ij}+\frac{1}{4}\tilde{k}^{2}.$ (10)
@@
$\frac{d\tilde{k}}{dt}=\frac{d}{dt}\left(\tilde{b}^{ij}\tilde{k}_{ij}\right)=-\tilde{b}^{mi}\tilde{b}^{nj}\tilde{k}_{mn}\tilde{k}_{ij}+\tilde{b}^{ij}\frac{d}{dt}\tilde{k}_{ij},$ (11)
@@
$\tilde{R}=\frac{d\tilde{k}}{dt}+\frac{1}{4}\tilde{k}^{2}+\frac{1}{4}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{mn}\tilde{k}_{ij}.$ (12)
@@
$\tilde{G}_{00}=\tilde{R}_{00}-\frac{1}{2}\tilde{g}_{00}\tilde{R}=\frac{1}{8}\tilde{k}^{2}-\frac{1}{8}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{mn}\tilde{k}_{ij}.$ (13)
@@ 2.2.1 The radiation perfect fluid
The first kind of source, which is instructive but turns out to have no interesting solutions, is the radiation perfect fluid, for which the stress tensor is
@@ 2.2.2 The source-free magnetic field
$\tilde{\eta}_{ijk}=\tilde{a}^{1/2}\epsilon_{ijk},$ (22)
@@ 2.2.3 The kinetic part
$\tilde{T}^{\mbox{kt}}_{00}$ $=$ $\frac{1}{\sqrt{\tilde{a}}}\int f(\tilde{b}^{mn}p_{m}p_{n})^{1/2}d^{3}p,$ (35)
$\tilde{T}^{\mbox{kt}}_{0i}$ $=$ $-\frac{1}{\sqrt{\tilde{a}}}\int fp_{i}d^{3}p,$ (36)
$\tilde{T}^{\mbox{kt}}_{ij}$ $=$ $\frac{1}{\sqrt{\tilde{a}}}\int fp_{i}p_{j}(\tilde{b}^{mn}p_{m}p_{n})^{-1/2}d^{3}p,$ (37)
@@ 3.1 Behaviour of the metric and curvature
From (7)–(9) the physical Ricci tensor can be written
@@ 3.2 Behaviour of the source-terms
$\tilde{T}^{\mbox{kt}}_{00}$ $=$ $\tau^{-4}\frac{1}{\sqrt{a}}\int f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p,$ (56)
$\tilde{T}^{\mbox{kt}}_{0i}$ $=$ $-\tau^{-3}\frac{1}{\sqrt{a}}\int fp_{i}d^{3}p,$ (57)
$\tilde{T}^{\mbox{kt}}_{ij}$ $=$ $\tau^{-2}\frac{1}{\sqrt{a}}\int fp_{i}p_{j}(b^{mn}p_{m}p_{n})^{-1/2}d^{3}p.$ (58)
@@ 3.3 The rescaled Einstein equations
$\frac{2}{\tau}Z_{ij},\quad Z_{ij}=\tau 8\pi\tilde{T}_{ij}-\frac{1}{\tau}a_{ij},$ (60)
@@
$Z_{ij}=\frac{1}{\tau}\left[\frac{8\pi}{\sqrt{a}}\int fp_{i}p_{j}\frac{d^{3}p}{(b^{mn}p_{m}p_{n})^{1/2}}+\frac{2}{a}(a_{ij}a_{mn}-2a_{im}a_{jn})h^{m}h^{n}-a_{ij}\right],$ (61)
@@
$\frac{d}{d\tau}a=ka,\quad\frac{d}{d\tau}b^{ij}=-{b}^{im}{b}^{jn}{k}_{mn},$ (64)
@@ 3.4 The Einstein constraint equations
The momentum constraint equation is
$\tilde{G}_{0i}-8\pi\tilde{T}_{0i}=0,$
which is equivalent to
$\frac{1}{\sqrt{\tilde{a}}}\int fp_{i}d^{3}p=0,$ (69)
@@
$\frac{1}{\tau^{3}\sqrt{{a}}}\int fp_{i}d^{3}p=0.$ (70)
Thus we impose $\int fp_{i}d^{3}p=0$ as a condition on the initial $f$ and this continues to hold since the Boltzmann equation satisfies
$\int p_{\alpha}\frac{\partial f}{\partial t}d^{3}p=0,$ (71)
which expresses energy-momentum conservation in collisions. On the other hand we have the Hamiltonian constraint as
@@
$\iff$ $\tilde{H}=\frac{1}{8}\tilde{k}^{2}-\frac{1}{8}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{mn}\tilde{k}_{ij}-\Lambda-\frac{2}{\tilde{a}}\tilde{a}_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{\tilde{a}}}\int f(\tilde{b}^{mn}p_{m}p_{n})^{1/2}d^{3}p=0.$ (72)
@@
$\frac{d}{dt}\tilde{H}=$ $\frac{1}{4}\tilde{k}\frac{d\tilde{k}}{dt}+\frac{1}{4}\tilde{b}^{jn}\tilde{b}^{ip}\tilde{b}^{mq}\tilde{k}_{pq}\tilde{k}_{mn}\tilde{k}_{ij}-\frac{1}{4}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{mn}\frac{d}{dt}\tilde{k}_{ij}+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}$
$+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{b}^{mi}\tilde{b}^{nj}\tilde{k}_{ij}\tilde{T}_{mn}^{\mbox{kt}}$
$=$ $\frac{1}{2}\tilde{k}\left(-\frac{1}{4}\tilde{b}^{im}\tilde{b}^{jn}\tilde{k}_{ij}\tilde{k}_{mn}-\tilde{R}_{00}\right)+\frac{1}{2}\tilde{b}^{ip}\tilde{b}^{jq}\tilde{k}_{pq}\left(\frac{1}{4}\tilde{k}\tilde{k}_{ij}-\frac{1}{2}\tilde{k}_{im}\tilde{b}^{mn}\tilde{k}_{nj}-\tilde{R}_{ij}\right)$
$+\frac{1}{4}\tilde{b}^{jn}\tilde{b}^{ip}\tilde{b}^{mq}\tilde{k}_{pq}\tilde{k}_{mn}\tilde{k}_{ij}+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{b}^{mi}\tilde{b}^{nj}\tilde{k}_{ij}\tilde{T}_{mn}^{\mbox{kt}}$
$=$ $-\frac{1}{2}\tilde{k}\tilde{R}_{00}-\frac{1}{2}\tilde{k}^{ij}\tilde{R}_{ij}+8\pi\tilde{k}\tilde{T}^{\mbox{max}}_{00}-\frac{2}{\tilde{a}}\tilde{k}_{ij}h^{i}h^{j}+4\pi\tilde{k}\tilde{T}_{00}^{\mbox{kt}}+4\pi\tilde{b}^{mi}\tilde{b}^{nj}\tilde{k}_{ij}\tilde{T}_{mn}^{\mbox{kt}}$
$=$ $-\frac{1}{2}\tilde{k}\tilde{H},$ (73)
where we again used the energy-momentum conservation of collisions (71) to remove the term with $\frac{\partial f}{\partial t}$. We see that if the Hamiltonian constraint equation is satisfied initially it will always be satisfied.
@@
$H=\frac{1}{8}\left(\frac{k}{\tau}+\frac{6}{\tau^{2}}\right)^{2}-\frac{1}{8\tau^{4}}(\tau k_{ij}+2a_{ij})(\tau b^{mi}b^{nj}k_{mn}+2b^{ij})-\Lambda-\frac{2}{a\tau^{4}}a_{ij}h^{i}h^{j}$
$-\frac{8\pi}{\tau^{4}\sqrt{a}}\int f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p=0,$ (74)
@@
$\tau^{4}H=-\Lambda\tau^{4}+\frac{1}{8}\tau^{2}\left(k^{2}-k_{ij}b^{mi}b^{nj}k_{mn}\right)+k\tau+3-\frac{2}{a}a_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{a}}\int f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p=0.$ (75)
@@
$0=3-\frac{2}{a}a_{ij}h^{i}h^{j}-\frac{8\pi}{\sqrt{a}}\int f(b^{mn}p_{m}p_{n})^{1/2}d^{3}p.$ (76)
3.5 Main results
Now that all relevant equations have been introduced and denoting by $S_{2}({\mathbb{R}}^{3})$ the space of $3\times 3$ symmetric matrices, we are able to state our main theorem concerning solutions of the Einstein-Vlasov system with a magnetic field:
Theorem 1
$\tilde{a}_{ij}$ $=2ta_{0ij}+o(t^{\frac{3}{2}}),$ (77)
$\tilde{k}_{ij}$ $=2a_{0ij}+o(t^{\frac{1}{2}}).$ (78)
$\tilde{a}_{ij}$ $=2ta_{0ij}+o(t^{2}),$ (79)
$\tilde{k}_{ij}$ $=2a_{0ij}+o(t).$ (80)
The proof will be obtained using the conformally rescaled variables which will be done in the next section. Concerning solutions of the Einstein-Boltzmann system with a magnetic field our main result is as follows:
Theorem 2
with
with
$\mathcal{C}_{ij}=\frac{1}{(\gamma-1)^{2}}2^{\gamma}a_{2ij},$ (86)
where $a_{2ij}$ is a constant which depends only on $f_{0}$ and $h$.
As in the Vlasov case the proof will be obtained using the conformally rescaled variables which will be done Section 5.